OpenStudy (anonymous):

find the indicated function values of the function g(x)=6x-7 a.g(0) b.g(-9) c.g(-1/6) d.g(-9b) e.g(b+8)

jimthompson5910 (jim_thompson5910):

I'll do the first three to get you started a) g(x)=6x-7 g(0)=6(0)-7 g(0)=0-7 g(0)=-7 ------------------------------------------------------- b) g(x)=6x-7 g(-9)=6(-9)-7 g(-9)=-54-7 g(-9)=-61 ------------------------------------------------------- c) g(x)=6x-7 g(-1/6)=6(-1/6)-7 g(-1/6)=-6/6-7 g(-1/6)=-1-7 g(-1/6)=-8

jimthompson5910 (jim_thompson5910):

The basic method is to replace the variable with the given value and evaluating like normal.